5. Suppose a baseball player hits a single” twice as likely as a double”, which is twice as likely as a triple”, which is
Answer #1

Theory: - PDF , (Probability is a statistical experssim a probability distribution for random variable Density functions) tha| PB-1) + P(k-) + R (K=J) + R(B=4) = 0.3 P 15 10.075 이명 이 c. 031S R.D.F Ple) 0 10.15 1. 13 01-23 | 나. 0.01S OIS 0.037S TotalNow F C.D. Flx )= P(X£ x) P(B40) = P(8=0) = 0-7 PCBL 12 ple=0) +P18=3) = 0.7 + 0.15 = 0.85 p(342) = P(B-0)+ P(6-1) +P (639) .رود 5 . 76 ,3) او را ر (كلام . ( بن 2 , 00) ) . و 4 3 unة ، . No 2 1 م b ) - ) curve (...